# The Power of Few: Accelerating and Enhancing Data Reweighting with Coreset Selection

### v1

@@ 1 Introduction
In the evolving field of machine learning, numerous strategies have been developed to tackle the complexities of model training [1]. Two particularly noteworthy techniques are coreset selection[2, 3, 4, 4, 5, 6, 7, 8, 9, 10, 11], aimed at improving computational efficiency by selecting most representative training data or removing harmful training data, and data reweighting [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], designed for generalization or faster convergence. While each has contributed to advancements in machine learning, they come with their own sets of challenges. Coreset selection, for instance, can sometimes overlook the nuanced diversity inherent in larger datasets. On the other hand, data reweighting incurs a significant computational burden, particularly when applied to expansive datasets.

### v3

@@ 1 Introduction
In the evolving field of machine learning, numerous strategies have been developed to tackle the complexities of model training [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Two particularly noteworthy techniques are coreset selection[26, 27, 28, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], aimed at improving computational efficiency by selecting most representative training data or removing harmful training data, and data reweighting [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], designed for generalization or faster convergence. While each has contributed to advancements in machine learning, they come with their own sets of challenges. Coreset selection, for instance, can sometimes overlook the nuanced diversity inherent in larger datasets. On the other hand, data reweighting incurs a significant computational burden, particularly when applied to expansive datasets.